# On the Evaluation of Generative Models in Distributed Learning Tasks

### v3

@@ A.2 Proof of Theorem 2
Appendix B Details
We have trained WGAN-GP (Salimans et al., 2016) and DDPM (Ho et al., 2020) in a federated learning setting by utilizing FedAvg approach (McMahan et al., 2017). The experiment protocols for WGAN-GP and DDPM are copied from original works. The communication interval of FedAvg is set as 160 iterations for both WGAN-GP and DDPM. We have tried different communication intervals for both models. The communication frequency will affect model performance but have no influence on the conclusions in the main part of our paper.
Appendix C General 1-Wasserstein-Distance evaluation Metrics
@@
where $\Pi(\mathbb{P}_{g},\mathbb{P}_{t})$ denotes the set of all joint distribution $\lambda(x,y)$ whose marginal distribution are respectively $\mathbb{P}_{g}$ and $\mathbb{P}_{t}$. However, the direct estimation of $W(\mathbb{P}_{g},\mathbb{P}_{t})$ is highly intractable. On the other hand, the Kantorovich-Rubinstein duality (Villani et al., 2009) gives,
@@
Appendix D Evaluation of Synthetic Gaussian Mixture Data with the Log-Likelihood Score
We also evaluated the synthetic Gaussian mixture dataset mentioned in Section 5.1 with the standard log-likelihood (LL) score. In this experiment, we note that we have access to the probability density functions (PDF) of the simulated generator. We utilized the generator $G_{var_{x}}$ described in the main text and performed the evaluation over the parameter $\mathrm{var}_{x}$ in the range $[0,40]$. As can be shown in the general case, LL-avg and LL-all led to the same value for every evaluated model. As shown in Figure 9a, they reached their maximum value at $\mathrm{var}_{x}=2$. On the other hand, we set a new generator $G_{mean_{x}}$ generating samples according to $\mathcal{N}([\mathrm{mean}_{x},0]^{\top},\bm{\Sigma})$, where $\bm{\Sigma}=\text{diag}([2,1]^{T})$. We gradually increased $\mathrm{mean}_{x}$ from -2 to 2 and plotted LL-avg, LL-all, and LL-ref in Figure 9b.
(a) $G_{var_{x}}$
(b) $G_{mean_{x}}$
Appendix E Extra Experiment Results on Well-trained Generative Models
E.1 Extra samples from clients
@@
E.2 Results on AFHQ dataset
Utilizing a another model weight that is pre-trained on AFHQ-wild dataset (Choi et al., 2020), we extend the experiment described in Section 5.3. Similar to the main text, we gradually increase the truncation parameter of generators We show some examples from generators, clients on Figure 12 and Figure 11 and plot the relationship between FID-scores and KID-scores on Figure 13. The results that are illustrated in the figures still indicates that FID-avg can lead to different ranking with FID-all, while KID-avg can be treated as a more stable evaluation metric in the distributed learning setting.
@@ E.2 Results on AFHQ dataset
Appendix F Extra Experiment Results on Federated Datasets
Similar to the federated CIFAR10 in Section 5.2, federated CIFAR100 and federated ImageNet-32 are conducted by grouping samples from a each class. In this section, we report some extra numerical results that evaluated on the three federated image datasets.
F.1 Evaluate Sequence of Net-based Generator on Federated CIFAR10
@@
We trained the WGAN-GP(Gulrajani et al., 2017) generative models multiple times using different random states, and we set different training lengths for every training procedure. We saved the models at different checkpoints every $10$ epochs, which is common in training generative models to select the best-performing saved model according to an evaluation metric.
@@
F.2 Evaluate CIFAR100 Generator on Federated CIFAR10
@@
F.3 Evaluate CIFAR100 on Federated ImageNet-32
@@
Appendix G Extra Experiment Results on Variance-Limited Federated Dataset
@@
G.1 Results on Variance-Limited Federated CIFAR10
@@
G.2 Results on Variance-Limited Federated CIFAR100
@@
G.3 Results on Variance-Limited Federated IN32
@@
G.4 The Effect of Intra-Client Variance

### v4

@@ A.2 Proof of Theorem 2
A.3 Proof of Proposition 1
Consider the FID-all-minimizing parameters in Theorem 2 resulting in
$\mathrm{FID}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{\widehat{G}}\Bigr{)}\,$ $=\,\mathrm{FID}\bigl{(}\mathcal{N}(\widehat{\bm{\mu}},\widehat{C}),\mathcal{N}({\bm{\mu}_{\widehat{G}}},{C}_{\widehat{G}})\bigr{)}$
$=\,\bigl{\|}\widehat{\bm{\mu}}-\bm{\mu}_{\widehat{G}}\bigr{\|}^{2}_{2}+\mathrm{Tr}\Bigl{(}\widehat{C}+C_{\widehat{G}}-2\bigl{(}\widehat{C}C_{\widehat{G}}\bigr{)}^{1/2}\Bigr{)}.$
Note that since we assume the number of clients $k$ is less than the dimension of the embedding, there exists a unit-norm vector $\bm{\beta}$ ($\|\bm{\beta}\|_{2}=1$) in the embedding space that is orthogonal to all mean vectors $\bm{\mu}_{1},\ldots,\bm{\mu}_{k}$ and hence to their mean $\widehat{\bm{\mu}}=\frac{1}{k}\sum_{i=1}^{k}\bm{\mu}_{i}$. Given $u=\mathrm{Tr}\bigl{(}\sum_{i=1}^{k}\lambda_{i}\bigl{(}\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top}\bigr{)}\bigr{)}$, we then consider the generator $G^{\prime}$ with the following mean and covariance parameters:
$\bm{\mu}_{G^{\prime}}\,=\,\widehat{\bm{\mu}}+\sqrt{u}\bm{\beta},\qquad C_{G^{\prime}}=C_{\widehat{G}}-\sum_{i=1}^{k}\lambda_{i}\bigl{(}\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top}\bigl{)}\,=\,\sum_{i=1}\lambda_{i}C_{i}.$
We claim that the generators $\widehat{G}$ and $G^{\prime}$ lead to the same client-based FID scores as for every $i$
$\mathrm{FID}\bigl{(}P_{X_{i}},P_{G^{\prime}}\bigr{)}\,$ $=\,\mathrm{FID}\bigl{(}\mathcal{N}({\bm{\mu}}_{i},{C}_{i}),\mathcal{N}({\bm{\mu}_{{G}^{\prime}}},{C}_{{G}^{\prime}})\bigr{)}$
$=\,\bigl{\|}{\bm{\mu}}_{i}-\bm{\mu}_{{G}^{\prime}}\bigr{\|}^{2}_{2}+\mathrm{Tr}\Bigl{(}{C}_{i}+C_{{G}^{\prime}}-2\bigl{(}{C}_{i}C_{{G}^{\prime}}\bigr{)}^{1/2}\Bigr{)}$
$=\bigl{\|}{\bm{\mu}}_{i}-\bm{\mu}_{\widehat{G}}\bigr{\|}^{2}_{2}+u+\mathrm{Tr}\Bigl{(}{C}_{i}+C_{{G}^{\prime}}-2\bigl{(}{C}_{i}C_{\widehat{G}}\bigr{)}^{1/2}\Bigr{)}-u$
$=\bigl{\|}{\bm{\mu}}_{i}-\bm{\mu}_{\widehat{G}}\bigr{\|}^{2}_{2}+\mathrm{Tr}\Bigl{(}{C}_{i}+C_{{G}^{\prime}}-2\bigl{(}{C}_{i}C_{\widehat{G}}\bigr{)}^{1/2}\Bigr{)}$
$=\mathrm{FID}\bigl{(}P_{X_{i}},P_{\widehat{G}}\bigr{)}.$
On the other hand, for the FID-all of $G^{\prime}$ we have
$\mathrm{FID}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{{G}^{\prime}}\Bigr{)}\,$ $=\,\mathrm{FID}\bigl{(}\mathcal{N}(\widehat{\bm{\mu}},\widehat{C}),\mathcal{N}({\bm{\mu}_{{G}^{\prime}}},{C}_{{G}^{\prime}})\bigr{)}$
$=\,\bigl{\|}\widehat{\bm{\mu}}-\bm{\mu}_{{G}^{\prime}}\bigr{\|}^{2}_{2}+\mathrm{Tr}\Bigl{(}\widehat{C}+C_{{G}^{\prime}}-2\bigl{(}\widehat{C}C_{{G}^{\prime}}\bigr{)}^{1/2}\Bigr{)}$
$=\,\bigl{\|}\widehat{\bm{\mu}}-\bm{\mu}_{\widehat{G}}\bigr{\|}^{2}_{2}+u\mathrm{Tr}\Bigl{(}\widehat{C}+C_{\widehat{G}}-2\bigl{(}\widehat{C}C_{\widehat{G}}\bigr{)}^{1/2}\Bigr{)}+u$
$=\,\bigl{\|}\widehat{\bm{\mu}}-\bm{\mu}_{\widehat{G}}\bigr{\|}^{2}_{2}+\mathrm{Tr}\Bigl{(}\widehat{C}+C_{\widehat{G}}-2\bigl{(}\widehat{C}C_{\widehat{G}}\bigr{)}^{1/2}\Bigr{)}$
$\;\;+2\mathrm{Tr}\bigl{(}\sum_{i=1}^{k}\lambda_{i}\bigl{(}\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top}\bigr{)}\bigr{)}$
$=\,\mathrm{FID}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{\widehat{G}}\Bigr{)}$
$\;\;+2\mathrm{Tr}\bigl{(}\sum_{i=1}^{k}\lambda_{i}\bigl{(}\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top}\bigr{)}\bigr{)}$
Therefore, Proposition 1’s proof is complete.
Appendix B General 1-Wasserstein-Distance evaluation Metrics
@@
where $\Pi(\mathbb{P}_{g},\mathbb{P}_{t})$ denotes the set of all joint distribution $\lambda(x,y)$ whose marginal distribution are respectively $\mathbb{P}_{g}$ and $\mathbb{P}_{t}$. However, the direct estimation of $W(\mathbb{P}_{g},\mathbb{P}_{t})$ is highly intractable. On the other hand, the Kantorovich-Rubinstein duality [23] gives,
@@
Appendix C Extra Experiment Results on FFHQ and AFHQ
C.1 Extra samples from clients
@@
C.2 Results on AFHQ dataset
Utilizing a another model weight that is pre-trained on AFHQ-wild dataset [26], we extend the experiment described in Section 5.2. Similar to the main text, we gradually increase the truncation parameter of generators We show some examples from generators, clients on Figure 10 and Figure 9 and plot the relationship between FID-scores and KID-scores on Figure 11. The results that are illustrated in the figures still indicates that FID-avg can lead to different ranking with FID-all, while KID-avg can be treated as a more stable evaluation metric in the distributed learning setting.
@@ C.2 Results on AFHQ dataset
Appendix D Evaluation on Synthetic Gaussian Mixture Datasets
As discussed in main text, the optimal selection of the covariance matrix differs for the FID-all and FID-avg aggregate scores. To illustrate this distinction, we performed a toy experiment, revealing that FID-avg attains its minimum value when the generator’s variance closely approximates that of an individual client, whereas FID-all is minimized when the variance equals that of the aggregate distribution.
[…]
The Log-Likelihood Score. We also evaluated the synthetic Gaussian mixture dataset with the standard log-likelihood (LL) score. In this experiment, we note that we have access to the probability density functions (PDF) of the simulated generator. We utilized the generator $G_{var_{x}}$ described in the main text and performed the evaluation over the parameter $\mathrm{var}_{x}$ in the range $[0,40]$. As can be shown in the general case, LL-avg and LL-all led to the same value for every evaluated model. As shown in Figure 13a, they reached their maximum value at $\mathrm{var}_{x}=2$. On the other hand, we set a new generator $G_{mean_{x}}$ generating samples according to $\mathcal{N}([\mathrm{mean}_{x},0]^{\top},\bm{\Sigma})$, where $\bm{\Sigma}=\text{diag}([2,1]^{T})$. We gradually increased $\mathrm{mean}_{x}$ from -2 to 2 and plotted LL-avg, LL-all, and LL-ref in Figure 13b.
(a) $G_{var_{x}}$
(b) $G_{mean_{x}}$
Appendix E Evaluation on Different Federated Image Datasets
E.1 Experiment Settings
We evaluated our theoretical results on standard image datasets. In our experiments, we simulated heterogeneous federated learning experiments consisting of non-i.i.d. data at different clients: for CIFAR-10 [19], we considered 10 clients, each owning 5000 samples exclusively from a single class of the image dataset. Therefore, every client’s dataset contains 5000 images having the same label. Similar to the federated CIFAR10, federated CIFAR100 and federated ImageNet-32 are conducted by grouping samples from a each class.
Neural Net-based Generators. We have trained WGAN-GP [14] and DDPM [3] in a federated learning setting by utilizing FedAvg approach [6]. The experiment protocols for WGAN-GP and DDPM are copied from original works. The communication interval of FedAvg is set as 160 iterations for both WGAN-GP and DDPM. We have tried different communication intervals for both models. The communication frequency will affect model performance but have no influence on the conclusions in the main part of our paper.
E.2 FID-based and KID-based Evaluation of Generative Models.
We evaluated the generative models according to FID-all, FID-avg, KID-all, and KID-avg as defined in Section 4. In several cases, we observed that FID-all / FID-avg could assign inconsistent rankings to the generators. Specifically, we computed FID-all and FID-avg for the ideal ’airplane’-class-based generator and neural net-based DDPM generators under the distributed CIFAR10 setting. We present some examples generated from the two generators in Figure 14 and report their scores according to the four metrics. The results suggest that FID-avg assigns a considerably lower score to the ideal ’airplane’-based generator, whose images preserve perfect details but lack diversity in image categories. Conversely, FID-all assigns a relatively lower value to the DDPM model because its images possess greater diversity. On the other hand, we also observed that KID-avg and KID-all give consistent rankings. Both of them led to the evaluation that the ideal plane generator is slightly better than the DDPM generator. In our implementation of KID-scores, we utilized the standard implementation of KID measurement from data with a polynomial kernel, $k(\bf{x},\bf{y})=\left(\frac{1}{d}\bf{x}^{T}\bf{y}+1\right)^{3}$, where $d$ is the dimension of feature vector. We note that our theoretical finding on the evaluation consistency under KID-all and KID-avg applies to every kernel similarity function. We also test KID-scores with a Gaussian RBF kernel $k_{\sigma}^{\text{rbf}}(\bf{x},\bf{y})=\exp\left(-\frac{1}{2\sigma^{2}}\|\bf{x}-\bf{y}\|^{2}\right)$ as formulated in [5], where we chose $\sigma=\sqrt{d}$ in the experiments. For images generated by diffusion model $\text{KID}^{\text{rbf}}$-all gives $4.277e^{-3}$ while $\text{KID}^{\text{rbf}}$-avg gives $4.295e^{-3}$. And for the airplane images in CIFAR10, $\text{KID}^{\text{rbf}}$-all gives $4.283e^{-3}$ while $\text{KID}^{\text{rbf}}$-avg gives $4.301e^{-3}$. The results indicate that for Gaussian RBF kernel $k_{\sigma}^{\text{rbf}}$, $\text{KID}^{\text{rbf}}$-all and $\text{KID}^{\text{rbf}}$-avg still gives consistent results. In this case, the $\text{KID}^{\text{rbf}}$-based evaluation suggests the images sampled from the diffusion model have higher quality than the set of airplane images in the CIFAR10 dataset.
E.3 Evaluate Sequence of Net-based Generator on Federated CIFAR10
@@
We trained the WGAN-GP[27] generative models multiple times using different random states, and we set different training lengths for every training procedure. We saved the models at different checkpoints every $10$ epochs, which is common in training generative models to select the best-performing saved model according to an evaluation metric.
@@
E.4 Evaluate CIFAR100 Generator on Federated CIFAR10
@@
E.5 Evaluate CIFAR100 on Federated ImageNet-32
@@
Appendix F Evaluation on Variance-Limited Federated Datasets
F.1 Experiment Setting
@@
F.2 Results on Variance-Limited Federated CIFAR10
@@
F.3 Results on Variance-Limited Federated CIFAR100
@@
Figure 17: Evaluation on Variance-Limited Federated CIFAR10.
@@
F.4 Results on Variance-Limited Federated IN32
@@
F.5 The Effect of Intra-Client Variance
@@
F.6 Precision/Recall and Density/Coverage Evaluations
Alongside FID and KID metrics, we expanded our analysis to include Precision/Recall [16] and Density/Coverage [17], adhering to the experimental frameworks outlined in Equation 2 and Equation 1.
Experiment Setup. For the Precision/Recall assessments, we employed the official implementation, setting the number of clusters to five, and evaluated the precision and recall metrics on a distributed CIFAR-10 dataset. In the case of Density/Coverage, we used the publicly available version of the evaluation metrics.
Results. As depicted in Figure 22, the precision and recall evaluations revealed variability in the consistency of rankings among generative models within a heterogeneous data environment. Specifically, while recall measurements (Recall-all and Recall-avg) consistently yielded identical values, the precision scores varied. This inconsistency in precision scores highlights potential challenges in using precision to rank models under distributed data distributions. Regarding Density and Coverage, our findings indicate that although density-based metrics (Density-all and Density-avg) sometimes provided inconsistent rankings, coverage-based assessments paralleled the recall metrics in consistently ranking models across evaluations.
Figure 22: Precision / Recall and Density / Coverage on distributed CIFAR10.